OpenStudy (anonymous):

Trig: ODD AND EVEN FUNCTIONS If cos(x) = -1/3, find cos(-x) Why is the result -1/3?

OpenStudy (anonymous):

we know cos x is an even function

OpenStudy (anonymous):

that is f(x) = f(-x) so if cosx = 1/3 then cos(-x) = 1/3 also :)

OpenStudy (anonymous):

and it is -1/3 for cos x and cos -x because by the definition of an even funtion